# Detection and Geographic Localization of Natural Objects in the Wild: A Case Study on Palms

### v1

@@ 1 Introduction
Palms (family Arecaceae) are vital to tropical ecosystems, serving as essential resources for pollinators and frugivores and influencing the evolution of dependent fauna [eiserhardt2011geographical, zambrana2007diversity]. Beyond their ecological roles, palms are deeply integrated in the livelihoods of rural and indigenous communities, providing food, construction materials, fuel, and medicine while supporting sustainable non-timber forest product markets. Particularly in regions like the Amazon, palms support subsistence practices and enhance resilience to socio-economic and environmental changes, reflecting the intricate links between biodiversity, human well-being, and sustainability [pitman2014distribution, malhi2014tropical, van2019palm, terborgh1986community]. Their distinctive star-shaped crowns make them well-suited for automated mapping using UAV imagery, supporting key efforts in biodiversity monitoring [sutherland2013identification, wagner2020regional].
@@
Figure 1: Palm Distribution Comparison. The first three images from previous studies [gibril2021deep, jintasuttisak2022deep] feature evenly spaced palms or clear backgrounds, while the last represents our case with natural spacing, occlusions, and complex backgrounds in tropical forests.
@@
Orthomosaic images are a valuable tool in remote sensing, with applications across various fields such as construction site monitoring, agricultural planning, environmental impact assessments, property development, and surveying, where accurate spatial data is essential [kucharczyk2021remote]. These images are created by stitching together hundreds or thousands of geolocated aerial frames into spatially coherent maps, enabling consistent analysis across landscapes. By providing a top-down view of large geographical areas, orthomosaics allow for precise measurement and detailed analysis of land features. Unlike standard aerial imagery, which contains perspective distortions, orthomosaics offer an in-scale representation of the ground. Low-flying drones enable efficient data collection without the need for extensive ground surveying [olson2021review].
@@
However, orthomosaics are not error-free and can be challenging to process. They often contain noise, artifacts, and stitching errors due to limited camera perspectives, variable lighting, wind-induced movement, and dynamic cloud cover during image acquisition [zhang2023aerial]. Additionally, resolution loss is a frequent byproduct of the stitching process. Typically stored in GeoTIFF format, orthomosaics can be exceptionally large, particularly in biodiversity conservation, forest health, and sustainable management applications, where a single file may range from 1 to 200 GB. As a result, image analysis must be performed in small patches that fit within the memory constraints of processing hardware.
@@
Object-based detection techniques, widely used in remote sensing, are well-established in computer vision. However, detecting and localizing naturally occurring objects in orthomosaics is inherently challenging [tagle2019identifying, cui2024palmprobnet, qin2021ag, jintasuttisak2022deep]. Unlike structured plantations, where trees are arranged in grid patterns with ample spacing (see Figure 1), palms in tropical forests grow irregularly, often obscured by dense vegetation that distorts key diagnostic features. The distribution of palm species is highly imbalanced, with only a few dominant species present in sufficient abundance to train reliable detection models [li2016deep, freudenberg2019large]. UAV-based imagery also introduces substantial lighting variability, as shadows from thick canopies and shifting sun angles create significant fluctuations in brightness and contrast across detection targets, further complicating model training and evaluation [gibril2021deep]. These challenges are amplified by the lack of high-quality labeled datasets, whose creation often requires extensive fieldwork and expert annotation [muscarella2020global, hidalgo2022sustainable].
@@
Lastly, manipulation of geographical information data requires domain expertise that is not readily available to computational scientists. A flexible pipeline that can process spatially referenced survey imagery and provide precise geographic target coordinates would support valuable downstream applications, leveraging state-of-the-art machine learning methods. These applications include monitoring palm abundance and distribution over time and across vast areas, and developing adaptive and modular survey methods. Additionally, balancing accuracy with interpretability and calibration is essential for trustworthy ecological decision-making [selvaraju2017grad, kuzucu2025calibration]. This study therefore introduces a novel end-to-end pipeline that addresses a major need in environmental monitoring, is computationally efficient and reliable, and can support variable-resource processing in field applications.
@@ 2.1 Palm Detection and Localization
[…]
@@ 2.2 Object Detection and Zero-Shot Segmentation
Object detection, a core computer vision task, identifies and localizes objects via bounding boxes [zou2023object] and underpins advanced applications such as image segmentation and object tracking [wang2022sygnet, ma2024rethinking, li2024CPDR]. The field is dominated by methods using You Only Look Once (YOLO) and Detection Transformer (DETR).
The YOLO family frames detection as a regression task balancing speed and accuracy. These methods often generate overlapping detections, which are typically resolved by a handcrafted process known as non-maximum suppression (NMS). YOLOv8 [yolov8] enhances detection through advanced backbone and neck architectures for feature fusion, and an anchor-free detection head optimized for accuracy and speed. YOLOv9 [yolov9] introduces programmable gradient information and the generalized efficient layer aggregation network to address information loss. YOLOv10 [yolov10] eliminates NMS through consistent dual assignments during training and one-to-one inference matching, coupled with a refined CSPNet backbone and a lightweight classification head to reduce computational cost. YOLO11 [yolo11] further enhances performance with a refined CSP bottleneck, hybrid attention, and adaptive anchors with extended IoU loss.
DETR [carion2020end] directly predicts object sets using learned queries, bypassing the need for post-processing such as NMS. DINO [zhang2022dino] enhances DETR with contrastive denoising and hybrid query initialization, while DDQ-DETR [zhang2023dense] introduces dense query assignment for improved one-to-one inference matching. RT-DETR [zhao2024detrs] optimizes DETR for real-time use via a hybrid encoder and multi-scale feature fusion.
Segment Anything Models (SAMs) are advanced segmentation models capable of segmenting any object in images using prompts such as points, boxes, or text [kirillov2023segment, ravi2024sam, mobile_sam]. Trained on the SA-1B (1 billion masks, 11 million images), SAM enables zero-shot inference and often surpasses fine-tuned methods in accuracy and efficiency [kirillov2023segment]. Its architecture features a ViT for image encoding, a prompt encoder to process input prompts, and a mask decoder that fuses features from both to generate segmentation masks. SAM 2 [ravi2024sam], trained on the SA-V dataset (50.9k videos, 642.6k masks), enhances video segmentation and object tracking by refining multi-scale feature extraction. Mobile SAM [mobile_sam] optimizes SAM for mobile use by simplifying the image encoder and using decoupled distillation, enhancing speed without compromising segmentation quality.
@@ 3 PALMS Dataset
Data in this study (See Figure 2) come from western Ecuador’s Choco tropical forest, including Fundación para la Conservación de los Andes Tropicales Reserve and adjacent Reserva Ecológica Mache-Chindul park (FCAT; 0023’28” N, 7941’05” W), Jama-Coaque Ecological Reserve (0006’57” S, 8007’29” W), Canande Reserve (031’34” N 7912’47” W), and Tesoro Escondido Reserve (033’16” N 7910’31” W). FCAT is a high diversity humid tropical forest at elevation $\sim$500m, receiving $\sim$3000 mm yr-1 precipitation with persistent fog during drier period. Jama-Coaque ranges from the boundary of the tropical moist deciduous/tropical moist evergreen forest at the lower elevations ($\sim$1000 mm precipitation yr-1, $\sim$250 m asl) to fog-inundated wet evergreen forests above 580m to 800m. Canande (350–500 m elevation) and Tesoro Escondido ($\sim$200 m elevation) are lowland everwet Choco forests, both receiving 4000–5000 mm yr-1 precipitation with no month experiencing drought stress or precipitation below 100 mm. These forests host several palm species with exposed canopy crowns, including the economically important Iriartea deltoidea, Socratea exorrhiza, and Oenocarpus bataua, with lesser amounts of Attalea colenda and Astrocaryum standleyanum, and species composition varying across study sites [browne2016diversity, lueder2022functional].
@@
To enable fine-tuned palm detection and zero-shot segmentation, we curated a dataset of 1,500 images (800$\times$800 pixels) from two FCAT reserve sites, capturing varied quality and palm density typical of natural forests. Manual annotation of palm crowns and isolated leaves faced challenges from vegetation overlap and orthomosaic distortions. Three trained experts initially labeled bounding boxes, followed by iterative refinement using a YOLOv8 detector trained on these annotations to identify and correct labeling errors or omissions. For landscape-scale validation, a trained expert annotated 5,026 georeferenced palm centers on four orthomosaics using ArcGIS Pro 3.3.1, with predicted coordinates from PRISM to refine annotations. This hybrid human-model workflow addressed the inherent complexity of labeling in natural forest environments, ensuring robust annotations for both detection and geospatial validation.
@@ 4 PRISM Pipeline
We selected YOLOv10 [yolov10] for its speed and performance. YOLOv10 introduces a consistent dual assignment strategy for NMS-free training, which combines one-to-many and one-to-one label assignments for enriched supervision while eliminating post-processing NMS. This approach uses a unified matching metric:
@@
We chose SAM 2 [ravi2024sam] for zero-shot segmentation due to its superior segmentation quality and improved speed compared to its predecessor. SAM 2 uses a hierarchical image encoder for multi-scale feature extraction and an optimized architecture that reduces computational overhead while maintaining high precision. Its efficient prompt encoding and memory attention mechanism enable rapid mask refinement during inference. During inference, bounding boxes from the detection model undergo NMS to remove duplicates, as patches are cropped from the orthomosaic with a stride. The remaining bounding boxes and their surrounding image regions serve as input prompts for SAM 2 to generate segmentation masks. Final outputs include visualized bounding boxes and masks, alongside georeferenced coordinates derived from NMS-cleaned bounding box centers, which are subsequently used to quantify counting accuracy.
@@
Model GFLOPS $\downarrow$ Params (M) $\downarrow$ FPS $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ $\text{AP}_{50}$uparrow $\text{AP}_{75}$uparrow mAP $\uparrow$
@@
Calibration analysis ensures model trustworthiness by aligning predicted confidences with empirical IoU, critical for ecological monitoring where overconfident false positives could mislead conservation decisions. We evaluate four calibration methods to quantify and mitigate errors in palm detection [kuzucu2025calibration]: (1) Linear regression (LR) maps logits to probabilities via a fitted linear function. (2) Isotonic regression (IR) fits a monotonic function by solving: $\min_{\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{n}}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$, subject to $\hat{y}_{i}\leq\hat{y}_{j},\forall i<j.$ (3) Temperature scaling (TS) divides logits by a learned temperature parameter $T$ before applying a sigmoid. (4) Platt scaling (PS) applies logistic regression to map logits to probabilities as: $p=\frac{1}{1+\exp(-(a\cdot\text{logit}+b))}$.
@@
Saliency maps generated by Grad-CAM [selvaraju2017grad] enhance interpretability by visualizing how the model identifies palms, aiding ecological validation and debugging. For a class score $y$, Grad-CAM computes gradients of $y$ with respect to feature maps $A^{k}$, yielding weights $\alpha_{k}=\frac{1}{Z}\sum_{i,j}\frac{\partial y}{\partial A_{ij}^{k}}$, where $Z$ is the number of spatial positions. The heatmap: $L_{\text{map}}=\text{ReLU}\left(\sum_{k}\alpha_{k}A^{k}\right)$ is then generated to highlight regions influencing predictions, such as palm crowns or diagnostic leaf patterns. This transparency helps ecologists verify whether the model focuses on biologically meaningful features (e.g., crown shapes) rather than spurious correlations (e.g., shadows).
@@ 5.1 Experimental Setting
where $J=25$ denotes the number of confidence bins, $D$ and $D_{j}$ represent the total detections and detections in bin $j$, and $p_{j}$ and $\text{IoU}_{j}$ are the average confidence and IoU within bin $j$. For $\text{LaACE}_{0}$, $p_{i}$ and $\text{IoU}_{i}$ correspond to per-detection confidence and IoU [kuzucu2025calibration]. We evaluated the counting performance using retrieval ratios between predicted and GT centers within $d=5$ meters and report the median shift of successful matches to quantify spatial deviation. For real-time analysis, we assess the inference speed on raw images that were directly taken from UAVs.
@@ 5.2 Detection and Segmentation Performance
[…]

### v2

@@ 1 Introduction
Palms (family Arecaceae) are vital to tropical ecosystems, serving as essential resources for pollinators and frugivores and influencing the evolution of dependent fauna [1, 2]. Beyond their ecological roles, palms are deeply integrated in the livelihoods of rural and indigenous communities, providing food, construction materials, fuel, and medicine while supporting sustainable non-timber forest product markets. Particularly in regions like the Amazon, palms support subsistence practices and enhance resilience to socio-economic and environmental changes, reflecting the intricate links between biodiversity, human well-being, and sustainability [3, 4, 5, 6]. Their distinctive star-shaped crowns make them well-suited for automated mapping using UAV imagery, supporting key efforts in biodiversity monitoring [7, 8].
@@
Figure 1: Palm Distribution Comparison. The first three images from previous studies [9, 10] feature evenly spaced palms or clear backgrounds, while the last represents our case with natural spacing, occlusions, and complex backgrounds in tropical forests.
@@
Orthomosaic images are a valuable tool in remote sensing, with applications across various fields such as construction site monitoring, agricultural planning, environmental impact assessments, property development, and surveying, where accurate spatial data is essential [11]. These images are created by stitching together hundreds or thousands of geolocated aerial frames into spatially coherent maps, enabling consistent analysis across landscapes. By providing a top-down view of large geographical areas, orthomosaics allow for precise measurement and detailed analysis of land features. Unlike standard aerial imagery, which contains perspective distortions, orthomosaics offer an in-scale representation of the ground. Low-flying drones enable efficient data collection without the need for extensive ground surveying [12].
@@
However, orthomosaics are not error-free and can be challenging to process. They often contain noise, artifacts, and stitching errors due to limited camera perspectives, variable lighting, wind-induced movement, and dynamic cloud cover during image acquisition [13]. Additionally, resolution loss is a frequent byproduct of the stitching process. Typically stored in GeoTIFF format, orthomosaics can be exceptionally large, particularly in biodiversity conservation, forest health, and sustainable management applications, where a single file may range from 1 to 200 GB. As a result, image analysis must be performed in small patches that fit within the memory constraints of processing hardware.
@@
Object-based detection techniques, widely used in remote sensing, are well-established in computer vision. However, detecting and localizing naturally occurring objects in orthomosaics is inherently challenging [14, 15, 16, 10]. Unlike structured plantations, where trees are arranged in grid patterns with ample spacing (see Figure 1), palms in tropical forests grow irregularly, often obscured by dense vegetation that distorts key diagnostic features. The distribution of palm species is highly imbalanced, with only a few dominant species present in sufficient abundance to train reliable detection models [17, 18]. UAV-based imagery also introduces substantial lighting variability, as shadows from thick canopies and shifting sun angles create significant fluctuations in brightness and contrast across detection targets, further complicating model training and evaluation [9]. These challenges are amplified by the lack of high-quality labeled datasets, whose creation often requires extensive fieldwork and expert annotation [19, 20].
@@
Lastly, manipulation of geographical information data requires domain expertise that is not readily available to computational scientists. A flexible pipeline that can process spatially referenced survey imagery and provide precise geographic target coordinates would support valuable downstream applications, leveraging state-of-the-art machine learning methods. These applications include monitoring palm abundance and distribution over time and across vast areas, and developing adaptive and modular survey methods. Additionally, balancing accuracy with interpretability and calibration is essential for trustworthy ecological decision-making [21, 22]. This study therefore introduces a novel end-to-end pipeline that addresses a major need in environmental monitoring, is computationally efficient and reliable, and can support variable-resource processing in field applications.
@@ 2.1 Palm Detection and Localization
Advancements in UAV technology, image stitching, and machine learning have driven significant progress in palm detection, segmentation, and localization from orthomosaic imagery. However, most studies have focused on commercially valuable species, such as oil and date palms, given their economic importance  [17, 9, 23, 10, 24]. For instance, [17] [17] used CNNs with a sliding window for oil palm counting in Malaysia, while [9] [9] developed a U-Net variant for enhanced date palm segmentation in UAE. [23] [23] proposed a Faster R-CNN variant with refined feature extraction and a hybrid class-balanced loss to monitor individual oil palm growth. More recently, YOLO-based approaches have been adopted: [10] [10] applied YOLOv5 for detecting date palms from UAV imagery over UAE farmlands, while [24] [24] employed YOLOv3 to detect and count oil palm trees for sustainable agricultural monitoring in Indonesia.
In contrast, the detection and localization of naturally occurring palms in tropical forests is largely underexplored. [14] [14] pioneered palm crown identification using random forest, showing machine learning’s potential for individual palm counting. [25] [25] applied a fully convolutional neural network with morphological operations to refine palm species segmentation. [8] [8] leveraged U-Nets and very high-resolution (0.5 m) multispectral imagery from the GeoEye satellite to map canopy palms over a large region of the Amazon rainforest.
@@ 2.2 Object Detection and Zero-Shot Segmentation
Object detection, a core computer vision task, identifies and localizes objects via bounding boxes [26] and underpins advanced applications such as image segmentation and object tracking [27, 28, 29]. The field is dominated by methods using You Only Look Once (YOLO) and Detection Transformer (DETR).
The YOLO family frames detection as a regression task balancing speed and accuracy. These methods often generate overlapping detections, which are typically resolved by a handcrafted process known as non-maximum suppression (NMS). YOLOv8 [30] enhances detection through advanced backbone and neck architectures for feature fusion, and an anchor-free detection head optimized for accuracy and speed. YOLOv9 [31] introduces programmable gradient information and the generalized efficient layer aggregation network to address information loss. YOLOv10 [32] eliminates NMS through consistent dual assignments during training and one-to-one inference matching, coupled with a refined CSPNet backbone and a lightweight classification head to reduce computational cost. YOLO11 [33] further enhances performance with a refined CSP bottleneck, hybrid attention, and adaptive anchors with extended IoU loss.
DETR [34] directly predicts object sets using learned queries, bypassing the need for post-processing such as NMS. DINO [35] enhances DETR with contrastive denoising and hybrid query initialization, while DDQ-DETR [36] introduces dense query assignment for improved one-to-one inference matching. RT-DETR [37] optimizes DETR for real-time use via a hybrid encoder and multi-scale feature fusion.
Segment Anything Models (SAMs) are advanced segmentation models capable of segmenting any object in images using prompts such as points, boxes, or text [38, 39, 40]. Trained on the SA-1B (1 billion masks, 11 million images), SAM enables zero-shot inference and often surpasses fine-tuned methods in accuracy and efficiency [38]. Its architecture features a ViT for image encoding, a prompt encoder to process input prompts, and a mask decoder that fuses features from both to generate segmentation masks. SAM 2 [39], trained on the SA-V dataset (50.9k videos, 642.6k masks), enhances video segmentation and object tracking by refining multi-scale feature extraction. Mobile SAM [40] optimizes SAM for mobile use by simplifying the image encoder and using decoupled distillation, enhancing speed without compromising segmentation quality.
@@ 3 PALMS Dataset
Data in this study (See Figure 2) come from western Ecuador’s Choco tropical forest, including Fundación para la Conservación de los Andes Tropicales Reserve and adjacent Reserva Ecológica Mache-Chindul park (FCAT; 0023’28” N, 7941’05” W), Jama-Coaque Ecological Reserve (0006’57” S, 8007’29” W), Canande Reserve (031’34” N 7912’47” W), and Tesoro Escondido Reserve (033’16” N 7910’31” W). FCAT is a high diversity humid tropical forest at elevation $\sim$500m, receiving $\sim$3000 mm yr-1 precipitation with persistent fog during drier period. Jama-Coaque ranges from the boundary of the tropical moist deciduous/tropical moist evergreen forest at the lower elevations ($\sim$1000 mm precipitation yr-1, $\sim$250 m asl) to fog-inundated wet evergreen forests above 580m to 800m. Canande (350–500 m elevation) and Tesoro Escondido ($\sim$200 m elevation) are lowland everwet Choco forests, both receiving 4000–5000 mm yr-1 precipitation with no month experiencing drought stress or precipitation below 100 mm. These forests host several palm species with exposed canopy crowns, including the economically important Iriartea deltoidea, Socratea exorrhiza, and Oenocarpus bataua, with lesser amounts of Attalea colenda and Astrocaryum standleyanum, and species composition varying across study sites [41, 42].
@@
To enable fine-tuned palm detection and zero-shot segmentation, we curated a dataset of 1,500 images ($\times800$ pixels) from two FCAT reserve sites, capturing varied quality and palm density typical of natural forests. Manual annotation of palm crowns and isolated leaves faced challenges from vegetation overlap and orthomosaic distortions. Three trained experts initially labeled bounding boxes, followed by iterative refinement using a YOLOv8 detector trained on these annotations to identify and correct labeling errors or omissions. For landscape-scale validation, a trained expert annotated 5,026 georeferenced palm centers on four orthomosaics using ArcGIS Pro 3.3.1, with predicted coordinates from PRISM to refine annotations. This hybrid human-model workflow addressed the inherent complexity of labeling in natural forest environments, ensuring robust annotations for both detection and geospatial validation.
@@ 4 PRISM Pipeline
We selected YOLOv10 [32] for its speed and performance. YOLOv10 introduces a consistent dual assignment strategy for NMS-free training, which combines one-to-many and one-to-one label assignments for enriched supervision while eliminating post-processing NMS. This approach uses a unified matching metric:
@@
We chose SAM 2 [39] for zero-shot segmentation due to its superior segmentation quality and improved speed compared to its predecessor. SAM 2 uses a hierarchical image encoder for multi-scale feature extraction and an optimized architecture that reduces computational overhead while maintaining high precision. Its efficient prompt encoding and memory attention mechanism enable rapid mask refinement during inference. During inference, bounding boxes from the detection model undergo NMS to remove duplicates, as patches are cropped from the orthomosaic with a stride. The remaining bounding boxes and their surrounding image regions serve as input prompts for SAM 2 to generate segmentation masks. Final outputs include visualized bounding boxes and masks, alongside georeferenced coordinates derived from NMS-cleaned bounding box centers, which are subsequently used to quantify counting accuracy.
@@
Model GFLOPS $\downarrow$ Params (M) $\downarrow$ FPS $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ $\text{AP}_{50}$ $\uparrow$ $\text{AP}_{75}$ $\uparrow$ mAP $\uparrow$
@@
Calibration analysis ensures model trustworthiness by aligning predicted confidences with empirical IoU, critical for ecological monitoring where overconfident false positives could mislead conservation decisions. We evaluate four calibration methods to quantify and mitigate errors in palm detection [22]: (1) Linear regression (LR) maps logits to probabilities via a fitted linear function. (2) Isotonic regression (IR) fits a monotonic function by solving: $\min_{\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{n}}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$, subject to $\hat{y}_{i}\leq\hat{y}_{j},\forall i<j.$ (3) Temperature scaling (TS) divides logits by a learned temperature parameter $T$ before applying a sigmoid. (4) Platt scaling (PS) applies logistic regression to map logits to probabilities as: $p=\frac{1}{1+\exp(-(a\cdot\text{logit}+b))}$.
@@
Saliency maps generated by Grad-CAM [21] enhance interpretability by visualizing how the model identifies palms, aiding ecological validation and debugging. For a class score $y$, Grad-CAM computes gradients of $y$ with respect to feature maps $A^{k}$, yielding weights $\alpha_{k}=\frac{1}{Z}\sum_{i,j}\frac{\partial y}{\partial A_{ij}^{k}}$, where $Z$ is the number of spatial positions. The heatmap: $L_{\text{map}}=\text{ReLU}\left(\sum_{k}\alpha_{k}A^{k}\right)$ is then generated to highlight regions influencing predictions, such as palm crowns or diagnostic leaf patterns. This transparency helps ecologists verify whether the model focuses on biologically meaningful features (e.g., crown shapes) rather than spurious correlations (e.g., shadows).
@@ 5.1 Experimental Setting
where $J=25$ denotes the number of confidence bins, $D$ and $D_{j}$ represent the total detections and detections in bin $j$, and $p_{j}$ and $\text{IoU}_{j}$ are the average confidence and IoU within bin $j$. For $\text{LaACE}_{0}$, $p_{i}$ and $\text{IoU}_{i}$ correspond to per-detection confidence and IoU [22]. We evaluated the counting performance using retrieval ratios between predicted and GT centers within $d=5$ meters and report the median shift of successful matches to quantify spatial deviation. For real-time analysis, we assess the inference speed on raw images that were directly taken from UAVs.
@@ 5.2 Detection and Segmentation Performance
Table 1 compares the detection performance when training and testing on different geographical regions within the FCAT reserve, introducing natural distribution shifts. YOLOv10 demonstrates superior efficiency with 177.04 FPS and 31.6M parameters while achieving competitive accuracy, particularly in higher-quality metrics like $\text{AP}{75}$ (67.94%) and mAP (61.73%). It marginally outperforms other YOLO variants by 0.17% AP75 and 0.11% mAP. DINO’s lower AP50 suggests reduced detection capability under shifted distributions. DDQ attains the best recall (85.66%) and a solid $\text{AP}_{50}$, but its lower FPS limits real-time applicability. RT-DETR achieves the highest precision (88.69%) but low recall (75.98%), indicating numerous missed palms. These results position YOLOv10 as optimal for speed-critical deployments, with DDQ reserved for high-recall scenarios.